it takes a plane 5 hours to travel 3000 miles with a tailwind, but it take the same plane 6 hours to travel the same distance with a headwind. hat is the speed of the plane in still air and speed of the wind?
Answer by josmiceli(19441)

p = the speed of the plane in still air
w = the speed of the wind
p + w = 3k / 5
p - w = 3k/6
add them
2p = (15K + 18K) / 30
2p = 33k / 30
p = 11k / 20
p = .55K
p = 550 mi/hr answer
p + w = 3k / 5
550 + w = 3K / 5
w = 600 - 550
w = 50 mi/hr
check
p - w = 3k/6
550 - 50 = 500
500 = 500
OK